Facebook, Inc. had no debt on its balance sheet in​ 2014, but paid ​$22 billion in taxes. Suppose Facebook were to issue sufficient debt to reduce its taxes by

​$270 million per year permanently. Assume​ Facebook's marginal corporate tax rate is 36 % and its borrowing cost is4.1 %

a. If​ Facebook's investors do not pay personal taxes​ (because they hold their Facebook stock in​ tax-free retirement​ accounts), how much value would be created​ (what is the value of the tax​ shield)?

b. How does your answer change if instead you assume that​ Facebook's investors pay a 20% tax rate on income from equity and a 39.6%

tax rate on interest​ income?
Answer #1

Case 1 All equity

Currently Facebook has no debt all equity company.

Current taxes paid = $ 22000 million

Tax Rate = 36%

Tax paid = 36% of Earnings before tax

earning before tax = 22000/0.36 = $ 61111.11millions

Case 2 Equity - Debt

Tax saved due to Debt = $ 270 million (given in question)

Tax = 22000-270 = $ 21730 million

Earning before tax = 21730/0.36 = $ 60361.11 million

Interest Paid = tax saved due to interest payment of debt / tax rate = 270/0.36 = $ 750 million

All Equity Equity- Debt
EBIT 61111.11 61111.11
Interest Payment@4.1% 0 750
Pretax Income 61111.11 60361.11
Taxes @ 36% 22000 21730

A. As taxes reduced due to idebt is $ 270 million per year permanently - perpetual debt

Tax shield = Tax * Debt

Total debt = Interest Paid / Interest Rate = 750 / 0.041 = $ 18292.68 million

Value of Tax shield = 0.36*18292.68 = 6585.366